OpenStudy (anonymous):

wenty-five members of the eighth grade class at Park Center Middle School are going to a museum and then to lunch. Each student must pay an entrance fee to the museum and $7.25 for lunch. The total cost for the trip is $443.75. What is the entrance fee for one student? I got 17. 46 443.75-7.25 then divided by 25 am i right?

OpenStudy (anonymous):

No, I don't think that's quite right...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

you only subtracted the cost of one student's lunch from the total... there should have been 25 lunches, each costing $7.25

OpenStudy (anonymous):

no the enternce fee is 7.25 they wanna know how much is the lunch since in total they paid 443.75

OpenStudy (anonymous):

The problem says lunch is $7.25, and you need to find entrance fee. Another way to set it up could be: 443.75 = 25(7.25 + Single Entrance Fee) then divide 443.75 by 25 to get the total cost per student... then subtract the cost of one lunch to get the amount of one student's entrance fee.

OpenStudy (anonymous):

lololol never mind!

OpenStudy (anonymous):

Ok :) Hope you figured it out :)

OpenStudy (anonymous):

so its 443.75 minus by 7.25 and divded by 25?

OpenStudy (anonymous):

so 17.46

OpenStudy (anonymous):

I get: total cost per student = 443.75 / 25 = 17.75 (lunch + entrance fee) = 17.75 7.25 + entrance fee = 17.75 entrance fee = 10.50 per student

OpenStudy (anonymous):

because 443.75=25(x+7.25) = 443.75= 25x+7.25 -7.25 =436.50/25=17.46

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

i get it thannks :D